OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -3, -12, -48, -192, ...

OpenStudy (freckles):

have you found the common ratio?

OpenStudy (freckles):

http://openstudy.com/users/abbs__#/updates/5654d0a2e4b0959c2b13a479 I have discussed this question in much detail you never asked any questions

rebeccaxhawaii (rebeccaxhawaii):

well what do you see the numbers are doing each time ?

OpenStudy (anonymous):

sorry i did not see you commented on it

rebeccaxhawaii (rebeccaxhawaii):

its okay tell me what you think is happening ?

OpenStudy (anonymous):

and i see the numbers being multiplied by 4

OpenStudy (freckles):

right that is your common ratio r=4

rebeccaxhawaii (rebeccaxhawaii):

yes so how would you get the ninth number ?

OpenStudy (freckles):

so now just notice your first term in plug into the form I had for your nth term

OpenStudy (anonymous):

how do i figure out whats the ninth number?

rebeccaxhawaii (rebeccaxhawaii):

we first you can keep multiplying the numbers until you get there

rebeccaxhawaii (rebeccaxhawaii):

so 192x4=768 768x4=3072 3072x4=12288 12288x4=49152 49152x4= your answer

OpenStudy (anonymous):

oh so the ninth number will be the ninth number is 296,608

OpenStudy (bibby):

the formula for the nth term in a geometric series is \(a_n=a_1*r^{n-1}\)

rebeccaxhawaii (rebeccaxhawaii):

no retype that

OpenStudy (anonymous):

i meant 196,608 sorry

rebeccaxhawaii (rebeccaxhawaii):

yes

OpenStudy (freckles):

even if they asked for the ninth term that is incorrect what you guys have as the ninth term they are looking for the nth term which is what I have tried to express to you the other day @Abbs__ and what @bibby is now telling you

OpenStudy (bibby):

oh, I didn't read the directions, I thought we wanted the 9th term lol

rebeccaxhawaii (rebeccaxhawaii):

oh me too

OpenStudy (anonymous):

I am very confused now

OpenStudy (freckles):

you are just missing negative sign on your ninth term

OpenStudy (freckles):

you need to identify the common ratio r which is 4 and the first term a1 which is -3 and just plug into the nth term of the sequence which is: \[a_n=a_1 \cdot r^{n-1}\]

OpenStudy (anonymous):

so it will be an=-3x4^n-1 ?

OpenStudy (anonymous):

aka an=-12^n-1 ?

OpenStudy (freckles):

yes to the first answer and no the second thing you wrote is not equivalent if you notice that is what @rebeccaxhawaii did above: \[\text{ ninth term }=a_9= =a \cdot r \cdot r \cdot r \cdot r \cdot r\cdot r\cdot r \cdot r=a r^{8}=a r^{9-1}\]

OpenStudy (anonymous):

oh okay i get it, thankyou! :)

OpenStudy (freckles):

\[a_n=-3 \cdot 4^{n-1} \text{ which is correct }\]